OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). A. y = -2x - 4 B. y = -2x - 8 C. y = -2x + 18 D. y = -2x – 12

OpenStudy (anonymous):

y-y=m(x-x) y+8=-2(x-2) y+8=-2x+4 y=-2x-4 so...A

OpenStudy (skullpatrol):

Or from y = mx + b -8 = (-2)(2) + b -8 = (-4) + b -4 = b therefore y =(-2)x -4 so...A